​Sam's Cat Hotel operates 52 weeks per​ year and uses a continuous review inventory system. It purchases kitty litter for ​$10.75 per bag. The following information is available about these bags.

  • Demand = 85 bags/week
  • Order cost = $57/order
  • Annual holding cost = 26% of cost per bag
  • Desired cycle-service level = 90%
  • Lead-time = 2 weeks
  • Standard deviation of weekly demand = 18 bags
  1. What is the economic order quantity (EOQ) for kitty litter?
  2. What would be the average time between orders (in weeks)?
  3. What should the reorder point R be?
  4. Assume that the firm decides to change to the periodic review system to control the inventory for kitty litter. The time between reviews, P, is the average time between orders as calculated in part b. What should be the target inventory level T?
  5. For the most recent review, an inventory clerk checked the inventory of kitty litter and found 310 bags. There were no open orders or backorders at the time. Determine how many units should the firm order for the periodic review system.
Answer #1

Reorder point is the level of inventory which the firm holds in stock and when the inventory level reach this point, the firm must reorder the item. Economic Order Quantity refers to the number of unit that are added to the inventory which can minimize the total inventory cost. It maintain a balance between setup costs and carrying costs.

n = 52 weeks d = 85 bags week S = $57/order H = 26% of cast cast = $10.15 H.26 x 10.75 = $2.195/unit/year CSL - 90% Z-value C18) ROP = (XLT) + (z J LT od) = (85x2)+(1.28 √ 2 x ) 70 + 32.58 = 202.58 or 203 unite Answer HP = 5 weeks T= d CP+ LT) + (2 J